Joni received $24,000 as part of a settlement in a class action lawsuit. She invested some money at 10% and the rest at 9% simple interest rates. If her total annual income from these investments was $2,274, how much did she invest at each rate?
Answer by checkley77(12844)

.10x+.09(24,000-x)=2,274
.10x+2160-.09x=2,274
.01x=2,274-2,160
.01x=114
x=114/.01
x=11,400 amount invested @ 10%.
24,000-11,400=12,600 amount invested @ 9%.
Proof:
.10*11,400+.09*12,600=2,274
1140+1134=2,274
2,274=2,274